Find the speed of a transverse wave that is in a 3 m string and has a mass of 0.130 kg if it can withstand a tension of 500 N.


1
Expert's answer
2021-09-09T17:21:31-0400

The speed of the wave in a string is given as follows (see https://courses.lumenlearning.com/suny-osuniversityphysics/chapter/16-3-wave-speed-on-a-stretched-string/):


"v = \sqrt{\dfrac{F}{\mu}}"

where "F = 500N" is the tension in the string, and "\mu" is the linear mass density.

By definition, have:


"\mu = \dfrac{m}{L}"

where "m = 0.130kg" is the total mass of the string, and "L = 3m" is its total length.

Thus, obtain:


"v = \sqrt{\dfrac{FL}{m}}\\
v = \sqrt{\dfrac{500N\cdot 3m}{0.13kg}} \approx 107m/s"

Answer. 107m/s.